OpenStudy (anonymous):

The volume of a cylinder is 93 ft3. The volume of a cube is 48 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (anonymous):

Volume of given cylinder=93 ft^3 (Volume of a cylinder=pie*r^2*h) Volume of a cone= (pie*r^2*h)/3 Therefore, volume of a cylinder that fits inside the cylinder=93/3=31 ft^3 Volume of given cube=48 ft^3 (Volume of a cube=lbh) Volume of a pyramid=lbh/3 Therefore, volume of the pyramid that fits inside the cube=48/3=16 ft^3 Difference of volume of a cylinder that fits inside the cylinder and volume of the pyramid that fits inside the cube is 31-16=15 ft^3 :)